What is the sound intensity of a whisper at a distance of 3.0 m, in W/m^2?
Express your answer in watts per meter squared. W/m2
Answer #1

Standard value of Intensity of whisper at 1.0 m is 1.0*10^-10 W/m^2

(I used this value from a reference book, If you have different value in your book let me know.)

Now intensity is given by:

I = Power/Area = P/A

I = P/(4*pi*R^2)

Since Power of whisper is constant, So Intensity will be inversely proportional to distance from source.

I2/I1 = (r1/r2)^2

r1 = 1.0 m

r2 = 3.0 m

I1 = 1.0*10^-10 W/m^2

So,

I2 = I1*(r1/r2)^2

I2 = 1.0*10^-10*(1.0/3.0)^2

I2 = 1.1*10^-11 W/m^2

Please Upvote.